OpenStudy (anonymous):

is 6x^2+5x-5 factorable?

OpenStudy (anonymous):

I came across this and I need to know if it can be factored, or if I need to do other ways of simplifying it

OpenStudy (jhannybean):

It's not easily factorable. T simplify something like this you either need to use the quadratic formula or the completing the square method.

OpenStudy (anonymous):

I also cam across this q10, it is quite messy but quadratic formula should do the trick

OpenStudy (jhannybean):

\[6x^2+5x-5\]\[x^2 +\frac{5}{6}x -\frac{5}{6}\]\[\left(x^2 +\frac{5}{6}x\right) -\frac{5}{6}\]\[\left(x^2 +\frac{5}{6}x+\frac{25}{144}\right)-\frac{5}{6} -\frac{25}{144}\]\[\left(x+\frac{5}{12}\right)^2-\frac{145}{144}\]

OpenStudy (jhannybean):

That would be your factored form in completing the square method (what I love to do)